# Observational predictions of some inflationary models

### v3

@@ 6 Conclusions
For the Starobinsky inflationary model, the inflationary potential is given by Eq. (14), and the slow–roll parameters are calculated analytically. The CMB angular power spectrum for this model is presented and compared with Planck $2018$ data in various plots, showing good agreement at small angular scales but with larger errors at intermediate scales. The acoustic peaks and troughs are also listed, and cosmic parameters calculated from the model are compared with observational values, demonstrating overall agreement.
The generalized Starobinsky inflationary model, described by Eq. (19), introduces an additional parameter $p$, and the slow–roll parameters are derived. The CMB angular power spectrum for various values of $p$ is presented, and the optimal value that best fits the Planck $2018$ data is determined to be $p=1.0004\pm 0.0001$. The model exhibits improved performance compared to the original Starobinsky inflationary model, with smaller errors in the relative error plots. The acoustic peaks and cosmic parameters are also analyzed, showing good agreement with observations.
Finally, the chaotic inflationary model with a step, described by Eq. (25), is investigated. The slow–roll parameters are derived, and the CMB angular power spectrum is presented. The model introduces a step in the potential, and its impact on the power spectrum is analyzed. The model is characterized by parameters such as $m$, $c$, $d$, and $\phi_{\textnormal{step}}$.
Overall, all three inflationary models successfully replicate the shape of the temperature power spectrum of the CMB while aligning with the values of cosmic parameters. However, the chaotic inflationary model with a step produce better results in the positioning and amplitudes of the peaks and troughs. The article concludes by summarizing the key findings and comparing the performance of the three models in reproducing the observed CMB angular spectrum, highlighting their strengths and limitations.
7 Acknolegments
We want to thanks to Professor Antony Lewis for your help always answering my question in his CosmoCoffee forum, and to Professor Victor Miralles for useful discussions about the use of CAMB code.

### v4

@@ 9 Conclusions
For the Starobinsky inflationary model, the inflationary potential is given by Eq. (15), and the slow–roll parameters are calculated analytically. The CMB angular power spectrum for this model is presented and compared with Planck $2018$ data in various plots, showing good agreement at small angular scales but with larger errors at intermediate scales. The acoustic peaks and troughs are also listed, and cosmological parameters calculated from the model are compared with observational values, demonstrating overall agreement.
The generalized Starobinsky inflationary model, described by Eq. (23), introduces an additional parameter $p$, and the slow–roll parameters are derived. The CMB angular power spectrum for various values of $p$ is presented, and the optimal value that best fits the Planck $2018$ data is determined to be $p=1.0004\pm 0.0001$. The model exhibits improved performance compared to the original Starobinsky inflationary model, with smaller errors in the relative error plots. The acoustic peaks and cosmological parameters are also analyzed, showing good agreement with observations.
While the generalized Starobinsky and chaotic models provide a much better fit to the CMB data than the original Starobinsky model, their low $\chi^{2}_{\nu}$ values suggest potential overfitting. Further investigations are necessary to evaluate whether these models’ complexity can be justified or if additional constraints are needed to prevent overfitting. Future work will focus on refining these models and exploring alternative inflationary scenarios that strike a balance between fit quality and model simplicity.
Regarding the reheating constraints of our inflationary models, the generalized Starobinsky inflationary model with $0.995\leq p\leq 1.0004$ and $\bar{\omega}_{\text{reh}}=-0.3$ offers a better positioning within the $(n_{\mathrm{S}},r)$ contour plot.
Finally, the chaotic inflationary model with a step, described by Eq. (30), is investigated. The slow–roll parameters are derived, and the CMB angular power spectrum is presented. The model introduces a step in the potential and its impact on the power spectrum is analyzed. The model is characterized by parameters such as $m$, $c$, $d$, and $\phi_{\textnormal{step}}$.
Overall, all three inflationary models successfully replicate the shape of the temperature power spectrum of the CMB while aligning with the values of cosmological parameters. The chaotic inflationary model with a step produce better results in the positioning and amplitudes of the peaks and troughs which could be verified by the value of $\chi^{2}$ calculated in Table 9. However, when performing a Bayesian analysis that considers variations in the effective number of neutrino species $N_{\text{eff}}$ across our three best inflationary models, the generalized Starobinsky inflationary model emerges as the favored model.
The article concludes by summarizing the key findings and comparing the performance of the three models in reproducing the observed CMB angular spectrum, highlighting their strengths and limitations.
10 Acknowledgements
We want to thank Professor Antony Lewis for his continuous support and assistance, always answering our questions on the CosmoCoffee Forum, and for his suggestion of using the Cobaya code. Additionally, we want to express our gratitude to Professor Víctor Miralles for the valuable discussions regarding the use of CAMB code, and Professor Rafael Hernández-Jiménez for his valuable explanations about the use of Cobaya code. We also greatly appreciate the suggestions made by the anonymous referees, which have helped us improve our manuscript.